A perfect rectangular prism is on top of a beam. The beam has two supports on each corner. The cross section of the beam is a circle (d = 7.05 cm).

Solid Steel box: length = 1.45 m , weight = 9000 N.

Length of beam = 1.45 m

What is the max normal compressive stress on the beam?


1
Expert's answer
2021-03-11T08:07:06-0500

S = "\pi" r3/4 = "\pi" d3/32 = 3.14 x (0.0705 m)3/32 = 3.44 x 10-5 m3

"\sigma" = My/I = 900 N x 1.45 m x 1.45 m / (3.44 x 10-5 m3 x 1.45 m) = 37.9 x 106 N/m2 = 38.0 MPa